OpenStudy (shadyykatie):

4.) The measures of the angles of a triangle are 2x, x + 24, and x – 4. Find the value of x and then find the measures of each of the angles. (Be sure to show all your work.)

OpenStudy (anonymous):

\(2x+x+24+x-4=180\) solve for \(x\)

OpenStudy (shadyykatie):

i am soo stuck on what to do next! You told me to sole for x but i don't know how

OpenStudy (anonymous):

\[2x+x+24+x-4=180\] combine like terms first, get \[4x+20=180\] subtract \(20\) from both sides get \[4x=160\] divide both sides by 4 and get \[x=40\]